OpenStudy (anonymous):

Find the terminal point determined by t= 10pi/3. I somewhat understand how to get the answer, but any extra help would be greatly appreciated.

jimthompson5910 (jim_thompson5910):

since this is on the unit circle, this means we can find any point of the form (x,y) where x = cos(t) y = sin(t) where t is the angle

jimthompson5910 (jim_thompson5910):

for example, if the angle was pi/3, then x = cos(t) x = cos(pi/3) x = 1/2 y = sin(t) y = sin(pi/3) y = sqrt(3)/2 This means that the terminal point for t = pi/3 is (1/2, sqrt(3)/2)

OpenStudy (anonymous):

Alright, so because I have 10pi/3, I would us pi/3 for the points?

jimthompson5910 (jim_thompson5910):

no pi/3 was just a similar example

jimthompson5910 (jim_thompson5910):

you would need to evaluate the cosine of 10pi/3 to get the x coordinate of the terminal point

jimthompson5910 (jim_thompson5910):

and you would need to evaluate the sine of 10pi/3 to get the y coordinate of the terminal point

jimthompson5910 (jim_thompson5910):

make sense?

jimthompson5910 (jim_thompson5910):

btw 10pi/3 is over 2pi, so you can subtract 2pi from it to get 10pi/3 - 2pi 10pi/3 - 6pi/3 (10pi - 6pi)/3 4pi/3

jimthompson5910 (jim_thompson5910):

this means that the angles 10pi/3 and 4pi/3 are coterminal angles

OpenStudy (anonymous):

I'm still pretty confused about it. Why is 10pi/3 over 2 pi?

jimthompson5910 (jim_thompson5910):

well I used a calculator to get 10pi/3 = 10.471975511966 and noticed it's larger than 6.28 (which is what 2pi is roughly equal to)

OpenStudy (anonymous):

Okay then why would I subract 2pi?

jimthompson5910 (jim_thompson5910):

2pi radians is the same as 360 degrees

jimthompson5910 (jim_thompson5910):

so if you add or subtract 2pi radians, you'll do a full 360 degree revolution and go from one angle and land on itself more or less

jimthompson5910 (jim_thompson5910):

ex: start at 0 degrees...add on 360 to get 0+360 = 360 degrees and you'll end up at the same spot

jimthompson5910 (jim_thompson5910):

so that's is why we can say that 10pi/3 and 4pi/3 are effectively the same angle

OpenStudy (anonymous):

Oh! So if instead of 10pi/3, I had something like 14pi/7, I would be able to subtract 2pi toget the terminal points?

jimthompson5910 (jim_thompson5910):

14pi/7 reduces to 2pi

jimthompson5910 (jim_thompson5910):

and you could leave it as 2pi or you can subtract 2pi from it to get 2pi - 2pi = 0pi = 0 this shows us that 0 and 2pi are coterminal points (ie effectively they are the same angle)

OpenStudy (anonymous):

@jim_thompson5910 , you are an absolute angel. Thank you so much!

jimthompson5910 (jim_thompson5910):

sure that's not "angle" lol jk glad I was of help though

jimthompson5910 (jim_thompson5910):

anyways, though, this may be very helpful http://www.regentsprep.org/Regents/math/algtrig/ATT5/600px-Unit_circle_angles_svg.jpg

jimthompson5910 (jim_thompson5910):

that's the unit circle and you'll see at 4pi/3 the point ( -1/2, -sqrt(3)/2) ), which is the terminal point for 10pi/3 (and 4pi/3)

OpenStudy (anonymous):

Thank you again :)

jimthompson5910 (jim_thompson5910):

sure thing